27/05/24
The carme-gsi computer has been moved from the server room into the shifter control room.
Startup of the DAQ looks OK. Settings loaded 2024Apr24
System wide checks OK. Aida16 clock status fails. Aida16 ADC calibration fails. Rest OK.
All FEEs look OK, we again see high rates in aida15 ~ 12k data items/sec, rate in other FEEs is 0.
Testing of the pulser widths has been done. Pulser widths will be uploaded soon.
Background run started. No beam, no target. File G22-201 and R2. Data being written is ~128 kb/s. Run will be left overnight, no night shift.
28/05/24
DAQ continues with alpha background run.
System wide checks OK. Aida16 clock status fails. Aida16 ADC calibration fails. Rest OK. Temps OK
Data can be seen in the spectrum browser on MIDAS. CARME monitoring code was started incorrectly yesterday. It has now be restarted and a small number of events are seen |